# the Morgan-Shalen compactification

### v1

@@ Abstract.
This paper studies the relationship between an analytic compactification of the moduli space of flat $\mathrm{SL}_{2}(\mathbb{C})$ connections on a closed, oriented 3-manifold $M$ defined by Taubes, and the Morgan–Shalen compactification of the $\mathrm{SL}_{2}(\mathbb{C})$ character variety of the fundamental group of $M$. We exhibit an explicit correspondence between $\mathbb{Z}/2$ harmonic 1-forms, measured foliations, and equivariant harmonic maps to $\mathbb{R}$-trees, as initially proposed by Taubes.
@@
$\mathbb{Z}$/2 harmonic spinors, $\mathbb{R}$-trees, Morgan-Shalen compactification
@@ 1. Introduction
The above results establish a bridge between the analytic and algebraic compactifications of the moduli space of flat $\mathrm{SL}_{2}(\mathbb{C})$ connections on $3$–manifolds. This allows us to prove new results on one side using results from the other side. We discuss several such applications in Section 7. Using the regularity results of Parker [Par23a], we prove a stronger regularity result for the Morgan–Shalen convergence (Corollary 7.7). Using Theorem 1.2, we show that the $\mathbb{Z}/2$ harmonic forms that arise as limits in Taubes’ construction must satisfy certain additional properties (Corollary 7.8). The Hubbard–Masur map [HM79] takes the space of measured foliations on a Riemann surface, modulo certain equivalence relations, to the space of quadratic differentials. Wolf [Wol95] showed that this map is a homeomorphism. Our construction of $\mathcal{H}$ can be interpreted as a generalization of the Hubbard–Masur map in dimension three, and it associates a canonical measured foliation with every boundary point in the Morgan–Shalen limit (see Section 7.4).
@@
Note that if a manifold has positive Ricci curvature, then the non-existence of $\mathbb{Z}/2$ harmonic $1$–forms follows directly from the Weitzenböck formula (see Equations (7), (8)).
@@
Condition (1) is a local regularity requirement for the zero locus of $\mathbf{v}$; it is weaker than the usual regularity assumption in the study of $\mathbb{Z}/2$ harmonic $1$–forms as stated [Don21, He23, Par23b]. The notations $\mu_{\mathbf{v}}$ and $d_{M,\mathbf{v}}$ in Condition (2) refer to the transverse invariant measure and the distance function on the leaf space defined by $\mathbf{v}$ (see (12) and (13)). We will show that Condition (2) always holds for $\mathbb{Z}/2$ harmonic $1$–forms that appear on the boundary of Taubes’ compactification (Corollary 7.8, part (2)).
It is worth noting that a completely different approach to Conjecture 1.3 is given in a forthcoming work of Parker [Par24], which uses the gluing constructions in gauge theory.
Acknowledgements. The authors wish to express their gratitude to many people for their interest and helpful comments. Among them are Nathan Dunfield, Xinghua Gao, Andriy Haydys, Zhenkun Li, Yi Liu, Ciprian Manolescu, Rafe Mazzeo, Tomasz Mrowka, Yi Ni, Jean-Pierre Otal, Greg Parker, Ao Sun, Clifford Taubes, and Mike Wolf.
@@ 6.4. Non-injectivity of the map $\overline{\Xi}$
Now by the product expression, $\hat{\rho}_{i}$ extend to representations $\rho_{i}$ of $\Gamma$, and they clearly converge to a Morgan–Shalen limit $\ell$, which is the extension of $\hat{\ell}$. Let $\widetilde{\Sigma}_{g}$ be the universal cover of $\Sigma_{g}$, and let $u:\widetilde{\Sigma}_{g}\to\mathcal{T}_{\ell}$ be an equivariant harmonic map with associated $\mathbb{Z}/2$ harmonic 1-form $\mathbf{v}$. We denote by $\mathcal{T}_{\mathbf{v}}$ the associated leaf tree defined by $\mathbf{v}$. On the other hand, each $\rho_{i}$ contains the free group $\mathbb{F}_{2}$ in its kernel, and therefore the edge stabilizers of $\mathcal{T}_{\ell}$ all contain free groups. By Skora’s Theorem [Sko96], $\mathcal{T}_{\ell}$ cannot be dual to a measured foliation on $\Sigma_{g}$. It follows that $\mathcal{T}_{\mathbf{v}}$ and $\mathcal{T}_{\ell}$ will necessarily have different length functions.
@@ 7. Applications
In this section, we give several applications of Theorems 1.1 and 1.2 and discuss some related results. In Section 7.1, we show that the projection map from a simply-connected manifold to the leaf space of a $\mathbb{Z}/2$ harmonic $1$–form may not be harmonic. In Section 7.2, we discuss a folklore conjecture on the non-existence of $\mathbb{Z}/2$ harmonic $1$–forms on $S^{3}$, and prove a result about $\mathbb{Z}/2$ harmonic forms on simply connected closed manifolds. In Section 7.3, we prove a $\mathcal{C}^{\infty}$ convergence result for Korevaar–Schoen limits using results from gauge theory. In Section 7.4, we discuss the relationship of Morgan–Shalen limits, singular measured foliations, and the Hubbard–Masur construction.
@@ 7.1. The quotient map to the leaf space
Since ${M}$ is compact, by the maximum principle for harmonic maps, every harmonic map from ${M}$ to an $\mathbb{R}$–tree must be constant. Therefore, the quotient map from ${M}$ to $\mathcal{T}_{{M},\mathbf{v}}$ cannot be harmonic.

### v2

@@ Abstract.
This paper studies the relationship between an analytic compactification of the moduli space of flat $\mathrm{SL}_{2}(\mathbb{C})$ connections on a closed, oriented 3-manifold $M$ defined by Taubes, and the Morgan–Shalen compactification of the $\mathrm{SL}_{2}(\mathbb{C})$ character variety of the fundamental group of $M$. We exhibit an explicit correspondence between $\mathbb{Z}/2$ harmonic 1-forms, measured foliations, and equivariant harmonic maps to $\mathbb{R}$-trees, as initially proposed by Taubes. As an application, we prove that $\mathbb{Z}/2$ harmonic 1-forms exist on all Haken manifolds with respect to all Riemannian metrics. We also show that there exist manifolds that support singular $\mathbb{Z}/2$ harmonic 1-forms but have compact $\mathrm{SL}_{2}(\mathbb{C})$ character varieties, which resolves a folklore conjecture.
@@
$\mathbb{Z}$/2 harmonic forms, $\mathbb{R}$-trees, Morgan-Shalen compactification
@@ 1. Introduction
The above results establish a bridge between the analytic and algebraic compactifications of the moduli space of flat $\mathrm{SL}_{2}(\mathbb{C})$ connections on $3$–manifolds. This allows us to prove new results on one side using results from the other side. We discuss several such applications in Section 7.
In Section 7.1, we prove an existence result of $\mathbb{Z}/2$ harmonic forms by considering harmonic maps to $\mathbb{R}$–trees.
A non-zero $\mathbb{Z}/2$ harmonic 1-form $\mathbf{v}$ is called non-trivial if $\mathbf{v}$ has non-trivial holonomy; in other words, if it is not given by a single-valued $1$–form. We say that $\mathbf{v}$ is singular if there exists a point on $M$ where $\mathbf{v}$ cannot be locally lifted to a single-valued form. Clearly, singular implies non-trivial, but the converse is false in general. Using the map $\mathcal{H}$ constructed in Section 6.1, we prove the following theorem.
Theorem 1.3.
Suppose $M$ is a rational homology sphere, and suppose there exists a closed connected embedded surface $S\subset M$ such that $S$ is two-sided, $\pi_{1}$–injective, and does not bound an embedded ball. Then for every Riemannian metric $g$ on $M$, there exists a non-trivial $\mathbb{Z}/2$ harmonic 1-form on $(M,g)$.
We remark that while all previous analytical constructions of non-trivial $\mathbb{Z}/2$ harmonic 1-forms and spinors require the metric to take specific forms ([DW21, TW20, TW24, He22, HP24, CH24]), Theorem 1.3 holds for all metrics on the given manifolds.
A rational homology sphere $M$ satisfies the conditions in Theorem 1.3 if and only if $M$ is reducible or Haken. On the other hand, if a closed oriented 3-manifold $M$ has $b_{1}(M)>0$, then there always exist trivial $\mathbb{Z}/2$ harmonic forms that are given by usual harmonic $1$–forms. Therefore, Theorem 1.3 has the following consequence.
Corollary 1.4.
Suppose a closed oriented 3-manifold $M$ is reducible or Haken. Then for every Riemannian metric $g$ on $M$, there exist $\mathbb{Z}/2$ harmonic 1-forms on $(M,g)$.
Since $\mathbb{Z}/2$ harmonic 1-forms describe the boundary of compactified moduli spaces of solutions to (1), it is natural to ask whether all of them can be deformed to solutions of the equation. The Kuranishi structure near the boundary of compactified moduli spaces for generalized Seiberg–Witten equations has been studied by Doan–Walpuski [DW20] and Parker [Par24b]. Equation (1) is a special case of the generalized Seiberg–Witten equations; in general, the compactified moduli spaces are described by $\mathbb{Z}/2$ harmonic spinors, which is a generalization of the concept of $\mathbb{Z}/2$ harmonic forms.
For the Seiberg–Witten equations with two spinors, Parker [Par24b] proved that all $\mathbb{Z}/2$ harmonic spinors, under some nondegenerate conditions, can be realized as limits of 1-parameter families of solutions to the equation. Motivated by Parker’s result, there has been a folklore conjecture that $\mathbb{Z}/2$ harmonic 1-forms could not exists on any 3-manifold with compact $\mathrm{SL}_{2}(\mathbb{C})$ character variety. A precise formulation of this conjecture was recently stated in [HP24, Conjecture 1.14].
Theorem 1.3 implies that there are counterexamples to this folklore conjecture. In fact, we have the following result.
Corollary 1.5.
There exist infinitely many closed 3-manifolds $M$ such that the $\mathrm{SL}_{2}(\mathbb{C})$ character variety of $\pi_{1}(M)$ is compact, and $M$ supports a singular $\mathbb{Z}/2$ harmonic 1-form with respect to every Riemannian metric.
Proof.
Boyer–Zhang [BZ98, Theorem 1.8] and Motegi [Mot88] constructed infinitely many closed oriented Haken 3-manifolds whose $\mathrm{SL}_{2}(\mathbb{C})$ character varieties are zero-dimensional. Since character varieties are affine varieties, zero-dimensional character varieties are compact. There are infinitely many examples $M$ in [Mot88] such that $H_{1}(M)$ are cyclic groups with odd orders. If $M$ is a rational homology sphere such that $H_{1}(M)$ has no 2-torsion, every $\mathbb{Z}/2$ harmonic form on $M$ is singular. Hence the result is proved by Theorem 1.3. ∎
Corollary 1.5 gives the first example of $\mathbb{Z}/2$ harmonic 1-forms (and more generally, $\mathbb{Z}/2$ harmonic spinors) that cannot be deformed into solutions of the corresponding gauge-theoretic equations. It also suggests that some properties of solutions to the Seiberg–Witten equations with two spinors, as studied in [DW20, DW21, Par24b], may not generalize to flat $\mathrm{SL}_{2}(\mathbb{C})$ connections.
@@
In Sections 7.2 to 7.5, we give several more applications of Theorems 1.1 and 1.2 and discuss some related results. Using the regularity results of Parker [Par23a], we prove a stronger regularity result for the Morgan–Shalen convergence (Corollary 7.8). Using Theorem 1.2, we show that the $\mathbb{Z}/2$ harmonic forms that arise as limits in Taubes’ construction must satisfy certain additional properties (Corollary 7.9). The Hubbard–Masur map [HM79] takes the space of measured foliations on a Riemann surface, modulo certain equivalence relations, to the space of quadratic differentials. Wolf [Wol95] showed that this map is a homeomorphism. Our construction of $\mathcal{H}$ can be interpreted as a generalization of the Hubbard–Masur map in dimension three, and it associates a canonical measured foliation with every boundary point in the Morgan–Shalen limit (see Section 7.5).
@@
Note that if a manifold has positive Ricci curvature, then the non-existence of $\mathbb{Z}/2$ harmonic $1$–forms follows directly from the Weitzenböck formula (see Eqns. (7), (8)).
@@
Condition (1) is a local regularity requirement for the zero locus of $\mathbf{v}$; it is weaker than the usual regularity assumption in the study of $\mathbb{Z}/2$ harmonic $1$–forms as stated [Don21, He23, Par23b]. The notations $\mu_{\mathbf{v}}$ and $d_{M,\mathbf{v}}$ in Condition (2) refer to the transverse invariant measure and the distance function on the leaf space defined by $\mathbf{v}$ (see (12) and (13)). We will show that Condition (2) always holds for $\mathbb{Z}/2$ harmonic $1$–forms that appear on the boundary of Taubes’ compactification (Corollary 7.9, part (2)). We note that a completely different approach to Conjecture 1.6 is given in a forthcoming work of Parker [Par24a], whch uses the gluing constructions in gauge theory.
Acknowledgements. The authors wish to express their gratitude to many people for their interest and helpful comments. Among them are Nathan Dunfield, Xinghua Gao, Andriy Haydys, Zhenkun Li, Yi Liu, Ciprian Manolescu, Rafe Mazzeo, Tomasz Mrowka, Yi Ni, Jean-Pierre Otal, Greg Parker, Ao Sun, Clifford Taubes, Thomas Walpuski and Mike Wolf.
@@ 6.4. Non-injectivity of the map $\overline{\Xi}$
Now by the product expression, the $\hat{\rho}_{i}$ extend to representations $\rho_{i}$ of $\Gamma$, and they clearly converge to a Morgan–Shalen limit $\ell$, which is the extension of $\hat{\ell}$. Let $\widetilde{\Sigma}_{g}$ be the universal cover of $\Sigma_{g}$, and let $u:\widetilde{\Sigma}_{g}\to\mathcal{T}_{\ell}$ be an equivariant harmonic map with associated $\mathbb{Z}/2$ harmonic 1-form $\mathbf{v}$. We denote by $\mathcal{T}_{\mathbf{v}}$ the associated leaf tree defined by $\mathbf{v}$. On the other hand, each $\rho_{i}$ contains the free group $\mathbb{F}_{2}$ in its kernel, and therefore the edge stabilizers of $\mathcal{T}_{\ell}$ all contain free groups. By Skora’s Theorem [Sko96], $\mathcal{T}_{\ell}$ cannot be dual to a measured foliation on $\Sigma_{g}$. It follows that $\mathcal{T}_{\mathbf{v}}$ and $\mathcal{T}_{\ell}$ will necessarily have different length functions.
@@ 7. Applications
In this section, we give several applications of Theorems 1.1 and 1.2 and discuss some related results.
Section 7.1 proves Theorem 1.3, which is an existence result for singular $\mathbb{Z}/2$ harmonic forms on a large class of rational homology spheres. In Section 7.2, we show that the projection map from a simply-connected manifold to the leaf space of a $\mathbb{Z}/2$ harmonic $1$–form may not be harmonic. In Section 7.3, we discuss a folklore conjecture on the non-existence of $\mathbb{Z}/2$ harmonic $1$–forms on $S^{3}$, and we prove a result about $\mathbb{Z}/2$ harmonic forms on simply connected closed manifolds. In Section 7.4, we prove a $\mathcal{C}^{\infty}$ convergence result for Korevaar–Schoen limits using results from gauge theory. In Section 7.5, we discuss the relationship of Morgan–Shalen limits, singular measured foliations, and the Hubbard–Masur construction.
7.1. Existence of singular $\mathbb{Z}/2$ harmonic 1-forms
We prove Theorem 1.3 using Culler–Shalen’s construction of dual trees associated with essential surfaces.
Let $M$ be a closed oriented 3-manifold with $\Gamma:=\pi_{1}(M)$. Assume $S\subset M$ is an embedded two-sided surface. We briefly review the construction of the dual tree of $S$ as in [Sha02, Sec. 1.4]. Let $p:\widetilde{M}\to M$ be the universal cover and $\widetilde{S}:=p^{-1}(S)$ the preimage of $S$. There is a canonical 1-dimensional simplicial complex $\mathcal{T}_{S}$ associated with $S$, defined as follows: the vertices are the connected components of $\widetilde{M}\setminus\widetilde{S}$, and the edges are the connected components of $\widetilde{S}$. An edge is incidental to a vertex if the corresponding component of $\widetilde{S}$ is contained in the boundary of the corresponding component of $\widetilde{M}\setminus\widetilde{S}$. The space $\mathcal{T}_{S}$ is a simplicial tree [Sha02, Sec. 1.4], and the fundamental group $\Gamma$ acts on $\mathcal{T}_{S}$ by simplicial homeomorphisms. In the following lemma, we show that under the assumptions in Theorem 1.3, the action of $\Gamma$ on $\mathcal{T}_{S}$ has no fixed points (see also [Sha02, Proposition 1.5.2]).
Lemma 7.1.
Suppose $M$ is a rational homology sphere. Suppose $S$ is a closed connected embedded surface in $M$ that is two-sided, $\pi_{1}$–injective, and does not bound an embedded ball in $M$. Then the action of $\Gamma$ on $\mathcal{T}_{S}$ has no global fixed points.
Proof.
Since $M$ is a rational sphere and $S$ is two-sided, $M\setminus S$ must be disconnected. In fact, if $M\setminus S$ is connected, then there is a circle in $M$ that intersects $S$ transversely at one point, which implies that the homology class of $S$ must be non-torsion, contradicting the assumptions. Since $S$ is connected, $M\setminus S$ has two connected components.
The action of $\Gamma$ on $\mathcal{T}_{S}$ has a globally fixed vertex if and only if there is a component $C$ of $M\setminus S$ such that $\pi_{1}(C)\to\pi_{1}(M)$ is surjective. By the Seifert–van Kampen theorem, this is equivalent to the surjectivity of $\pi_{1}(S)\to\pi_{1}(M\setminus C)$. Since $S$ is $\pi_{1}$–injective, this condition implies $\pi_{1}(S)\to\pi_{1}(M\setminus C)$ is an isomorphism. If $S$ is a sphere, then $M\setminus C$ must be a $3$–ball, which contradicts the assumptions. If the genus of $S$ is positive, by a standard result in 3-manifold topology (see, for example, [Hat07, Lemma 3.5]), the rank of the kernel of $H_{1}(S)\to H_{1}(M\setminus C)$ equals the genus of $S$, so $\pi_{1}(S)\to\pi_{1}(M\setminus C)$ cannot be isomorphic.
The action of $\Gamma$ on $\mathcal{T}_{S}$ has a globally fixed edge if and only if $\pi_{1}(S)\to\pi_{1}(M)$ is surjective. By the Seifert–van Kampen theorem and the assumption that $S$ is $\pi_{1}$–injective, this implies $\pi_{1}(S)\to\pi_{1}(M\setminus C)$ is an isomorphism for each component $C$ of $M\setminus S$. Hence we get a contradiction by the same argument as before. ∎
Proof of Theorem 1.3.
Let $\Gamma=\pi_{1}(M)$. Since $\mathcal{T}_{S}$ is a simplicial tree, it is complete as a metric space. By Lemma 7.1, the set of projective length functions $\mathcal{PL}(\Gamma)$ is non-empty. Hence the construction of the map $\mathcal{H}:\mathcal{PL}(\Gamma)\to\mathcal{M}_{\mathbb{Z}/2}$ in Section 6.1 implies that $\mathcal{M}_{\mathbb{Z}/2}$ is non-empty. Since $M$ is a rational homology sphere, every non-zero $\mathbb{Z}/2$ harmonic form on $M$ must be non-trivial. ∎
We remark that essential surfaces also played an important role in the recent construction of new Casson-type invariants by Dunfield and Rasmussen [DR24]; see also [DGR22] for results on the counting of essential surfaces.
@@ 7.2. The quotient map to the leaf space
Since ${M}$ is compact, by the maximum principle for harmonic maps, every harmonic map from ${M}$ to an $\mathbb{R}$–tree must be constant. Therefore, the quotient map from ${M}$ to $\mathcal{T}_{{M},\mathbf{v}}$ cannot be harmonic. ∎